OpenStudy (anonymous):

Find the degree measure of an arc that is 70% of its circle. Plz help will award medals

OpenStudy (anonymous):

@ganeshie8 plz help me

OpenStudy (solomonzelman):

 Like this ?

OpenStudy (solomonzelman):

Give me the radius and I'll find the angle.

OpenStudy (anonymous):

give me a second then @SolomonZelman

OpenStudy (anonymous):

i think it means that the arc covers 70% of the perimeter

OpenStudy (anonymous):

I do not have the radius and that is all the information i have

OpenStudy (anonymous):

yes yes you are correct @zupari

OpenStudy (solomonzelman):

Wait, so you just have an arc given? What about the area, do you know the area ? Or the diameter ?

OpenStudy (anonymous):

Find the degree measure of an arc that is 70% of its circle. Type in the number only. This is all i am given

OpenStudy (solomonzelman):

Anything that you know other than the `arc length` ?

OpenStudy (anonymous):

its just 360 * 70% = 252 degrees

OpenStudy (solomonzelman):

Noooo

OpenStudy (anonymous):

i know the measure of the arc is 70% of the circle

OpenStudy (anonymous):

wait is @zupari correct ??

OpenStudy (solomonzelman):

The formula says that \(\normalsize\color{blue}{ \rm{arc~~~length}=r\times θ }\) and thus, \(\normalsize\color{blue}{ \rm{arc~~~length} \div r= θ }\).

OpenStudy (solomonzelman):

But I only know the arc length.

OpenStudy (anonymous):

yes but if take it in degrees a circle covers 360 if its 70% then 252

OpenStudy (solomonzelman):

I think that if 70 is not units but percent, sorry for misreading the question-:( Then take 70 percent of 360. 360 × 0.70 = ?

OpenStudy (anonymous):

so the answer is 252 degrees

OpenStudy (anonymous):

THX guys